1. The Bradley Corporation produces a product with the following costs as of July 1, 20X1:

Material $3 per unit

Labor 3 per unit

Overhead 1 per unit

Beginning inventory at these costs on July 1 was 3,100 units. From July 1 to December 1, 20X1, Bradley Corporation produced 12,200 units. These units had a material cost of $2, labor of $3, and overhead of $4 per unit. Bradley uses LIFO inventory accounting.

a. Assuming that Bradley Corporation sold 13,400 units during the last six months of the year at $14 each, what is its gross profit?

Gross Profit ________

b. what is the value of ending inventory?

Ending Inventory ________

2. The Hartnett Corporation manufactures baseball bats with Pudge Rodriguez’s autograph stamped on them. Each bat sells for $49 and has a variable cost of $26. There are $37,950 in fixed costs involved in the production process.

a. Compute the break-even point in units.

break-even point ____________units

b. Find the sales (in units) needed to earn profit of $19,320.

Sales quantity needed _________units
Answer #1

1.

Inventory cost per unit of 3100 units = 3+3+1=       7  
Inventory cost per unit of 12200 units = 2+3+4=       9  


As per LIFO method, inventory that comes last is issued first. So closing inventory in hand are old or out of beginning inventories          
and Sales units are new or latest inventory.          


Sales 13400 units @$14=       187600  
_______________________________________
cost of goods sold           
Firstly, 12200 units will be issued@ 9 =       109800  
Balance @ $7 units (13400-12200) =1200 issued= 8400  
__________________________________________  
Total Cost of goods sold =       118200  
          
          
Gross profit = Sales-COGS          
187600   -118200 = 69400  
          
So, gross profit is    $69,400.00      


(b) Closing in inventory remaining under LIFO will be old units          
Closing inventory will be Opening + New inventories   -Sales      
3100+12200-13400=   1900      
Inventory value = 1900 units @ $7=   13300      
          
so, ending inventory value is    $13,300.00      

Note: Separate first question is Answered as per Chegg policy